La présentation est en train de télécharger. S'il vous plaît, attendez

La présentation est en train de télécharger. S'il vous plaît, attendez

SPS LS1 Planning meeting (N016) 03/07/13 David Mcfarlane

Présentations similaires


Présentation au sujet: "SPS LS1 Planning meeting (N016) 03/07/13 David Mcfarlane"— Transcription de la présentation:

1 SPS LS1 Planning meeting (N016) 03/07/13 David Mcfarlane

2 Topics Lifts and Monte-charges. Planning. False floors open in the BAs Works between CCC & BA3 (Surface) RTLs E.C.Rs. Access rights and safety equipment Important Information. A.O.B

3 Lift Maintenance June/July LocationWorksDateInterruption BA3 Safety works on Monte-Charge 3 rd /4 th July Monte-Charge will be OUT OF SERVICE BA6 Safety works on Monte-Charge 5 th /8 th July Monte-Charge will be OUT OF SERVICE

4 EN UndergroundSur BA1 UndergroundSur BA2 UndergroundSur BA3 UndergroundSur BA4 UndergroundSur BA5 UndergroundSur BA6 TI8TI2 TRANSFER TUNNELS UndergroundSur BA7 UndergroundSur BA80/TCC Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. General maintenance Cold check out & Start-up DSO tests/HV tests EPC Commissioning tests Vacuum Ions to NA61 General Maintenance General Maintenance General Maintenance General maintenance General Maintenance UA9 works RF upgrade works EPC works Civil Engineering RF upgrade works RF Upgrade works Survey Horizontal Vertical smoothing Cables Irradiated Cables Cables Re-route Cabling Cables EL Studies Prep work in LSS1 Re-install equipment removed for cable works Survey SPS LS1 schedule Beam on Cool down/RP survey and Magnet tests Protons to the North Area Protons to LHC Cool down NO ACCESS Cool down Limited ACCESS Magnets Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Fibres CV works CV studies CV Controls upgrade Vacuum TT10 Consolidation 18 KV Septa works Kicker Septa works Kicker TT20 pipework Magnets 18 KV Civil Engineering Magnets Ring-main Vacuum Crab Cavities EPC works Cooling tower 863 Ring-main Cables Fibres 18 KV Civil Engineering in TT10 Fibres CV maintenance Cables Kicker Cables RAL Septa replacement

5 False floors lifted in Surface BAs BAsCabling pulling startCable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/ /10/ /04/ /08/ /02/ /04/2013 BA206/01/201411/03/2014 BA3 BA3 (Fibre Optics) 27/03/ /04/ /05/ /05/2014 BA427/05/201312/07/2013 BA5 BA5 (Fibre Optics) 27/03/ /09/ /05/ /10/2013 BA6 BA6 (Fibre Optics) 30/09/ /5/ /10/ /05/2013 BA716/10/201307/11/2013 Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BAs that involve lifting of the false floors MUST be reported to the Coordination team If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

6 Works between BA3 and CCC Installation of new Transformers Creation of the new cable trench Works inside Building 874 (this includes works on the outside of the building.

7 New Cable Trench 2.5 to 3 weeks is needed to complete this work. (2 weeks at the beginning of October)

8 Installation of new Transformers Activités EPC concernés Installation des transformateurs SPS pour les Dipoles et Quadrupoles : 8 unités, Planning initial : Juillet-Aout-Septembre 2013, Besoins spécifiques : accès complet requis pour les grues 55T-220T et une remorque entre le BA3 et la CCC côté JURA et MONT-BLANC pendant toute la durée de linstallation (en rouge sur la carte).

9 Installation of Transformers - The installation of transformers quadrupoles will side with 220T crane JURA side MONT BLANC access is required in order to remove the transformers trailer no work (trench or other) should therefore begin before the transformers are all in place. - Pose Transformers Dipole must be done with the crane 55T side JURA - The goal for the implementation of all transformers is late July Given the nature of the work handling, no co-activity in this area should start before the end of installation of transformers, - Small layout work side JURA are required to allow access to the 220T crane: it is a prerequisite to start the installation of transformers. This request had been anticipated for the installation of transformers, however, the realization of this work now become urgent in view of the reduction of time allocated to the installation phase. This schedule organized in the interest of all will be finally approved in the coming days by Christophe Coupat is the EPC Coordinator for the installation work.

10 New RTLs Concernant lachat de RTL supplémentaires pour le SPS, il y a 2 possibilités : Soit on achète des chariots identiques à ceux en service actuellement. (TRANSLIFT) Soit on relance un appel doffre, du retour des utilisateurs SPS et des besoins réels du SPS. On a pas forcément besoin de véhicules pouvant transporter 2 personnes et pouvant tracter une remorque de 1T dans une pente de 11%. On pourrait compléter le parc actuel par des tracteurs de type Pefra ou Volk e-boy (pays dorigine DE), tracteur pour 1 personne avec une capacité de traction de 500Kg à plat, ce qui devrait être suffisant pour la majorité des utilisateurs du SPS. Ces tracteurs ont également une grande plateforme pour le transport du matériel, qui remplace dans bien des cas une remorque. Il sont également plus étroits et facilitent les croisements dans le tunnel SPS. Par contre, cest pour seulement pour le transport de 1 personne, tout comme les RTL Pefra du LHC. Ce type déquipement type e-boy, est à mon avis bien plus robuste, plus fiable et également moins chèr que les modèles Translift (env. 10%) Concernant le délai de livraison, cest des équipements non-standards, il faut compter env. 15 semaines

11 RTLs (The choice) TRANSLIFT PEFRA E-BOY

12 Engineering Change Requests Templates and Guidelines for Engineering Change Requests: – EDMS: SPS ECR template (in DFS): – https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx https://dfsweb.web.cern.ch/dfsweb/Services/DFS/DFSBrowser.aspx/A pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx

13 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 1.Removal of AEWB Pick-up from position to (EDMS No ) RELEASED 2.Removal of a BPWA Pick-up from position to (EDMS No ) RELEASED 3.Installation of a new Wideband Kickers system in position (EDMS No ) RELEASED 4.Installation of a new HBW Damper Pick-up in place of BPWA & AEWB (EDMS No ) RELEASED 5.Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No ) RELEASED 6.Reinstallation of pump and gauges (Pirani + Penning) in (vacuum sector 150) (EDMS No ) RELEASED 7.Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No v.2) RELEASED 8.Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No ) UNDER IMPLEMENTATION

14 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 9.Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No ) RELEASED 10.SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No ) RELEASED 11.Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No ) RELEASED 12.Installation d'un nouveau moniteur de position de type BPCR en LSS2 au (EDMS No ) RELEASED 13.Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No ) RELEASED 14.Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No ) RELEASED 15.Move of BPH to BPH (EDMS No ) RELEASED 16.Optical Fibre Racks. (EDMS No ) RELEASED

15 Current ECRs for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 17.Removal of old fixed TMP vacuum groups in (EDMS No ) RELEASED 18.BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). (EDMS No ) UNDER IMPLEMENTATION 19.Installation de Coffrets d'Alimentation en LSS1 et TT10. (EDMS No ) RELEASED 20.Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. (EDMS No ) RELEASED 21.Installation de nouveaux chemins de câbles en LSS1 et LSS2. (EDMS No ) UNDER APPROVAL

16 Access rights and safety equipment Everyone who enters the SPS MUST have the following: – Completed the correct courses. Safety course levels 1,2 and 3. Habilitation Electrique. Radiological Protection. Self Rescue Mask Training. – The correct access authorisation (via EDH). – A valid approved IMPACT. (with their name included on this IMPACT) – They must know the number of this IMPACT. – Safety shoes. – Biocell. – Electronic dosimetre. – Passive dosimetre. – Helmet plus Light. Please note that the medical service no longer contact you when your medical is about to expire!!!!!!! – You will receive an 2 months before informing you that some of your access rights will soon expire. – It is up to you to contact the medical service to arrange a medical.

17 Important information Useful Contact Numbers – SPS Coordination Team David MCFARLANE Franck BAIS – James RIDEWOOD – Control Room SPS CCC – TI CCC – CSA – RP Christophe Tromel – Angelito Herve – Nadine Conan (BA1) – Kurt Weiss (BA1) – Coordination Safety Emmanuel Paulat – John Etheridge

18 There are currently 39 IMPACTs that are LATE. If you have completed your work then you need to close the related IMPACT. A.O.B.

19 SPS coordination website \\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 17/07/ D20


Télécharger ppt "SPS LS1 Planning meeting (N016) 03/07/13 David Mcfarlane"

Présentations similaires


Annonces Google